OpenStudy (anonymous):

Use the T- table to determine 3 solutions to this equation:y=-4x-5 x y 0 1 -2

OpenStudy (campbell_st):

substitute x = 0 to find a y value y = 4 * 0 - 5 y = -5 repeat for x = 1 y = -4*(1) - 5 y = -9 and lastly x = -2 y = -4*(-2) - 5 y = 3 the table is